# HARMONIC: Cognitive and Control Collaboration in Human-Robotic Teams

### v2

@@ IV-B Development of Distributed Plans
Plans are distributed between strategic and tactical components in both UGV and drone systems, with each robot acting on reasoning from the OntoAgent in the strategic layer. The high-level plans (ontological script instances) are divided into steps communicated to the tactical layer as action commands, which are executed through low-level control encoded in BTs.
@@ IV-B1 High-level planning
The ontologies of agents in HARMONIC contain scripts for complex events that a particular agent can perform or understand other agents performing. Plans are instances of scripts with parameter values set for the situation at hand. These are instantiated by executing one or more of the ontologically stored meta-plans, that is, plans that HARMONIC robots instantiate to guide the generation of plans in their operational domain. One such meta-plan is the COLLABORATIVE-ACTIVITY plan, whose purpose is to help agents operating in teams organize themselves to accomplish a goal.
The details of the COLLABORATIVE-ACTIVITY plan vary based on the agent’s role in the team. For example, the team (or task) leader’s plan is typically more complex than the plans of its subordinates since team leaders select the plan to accomplish the target goal, resolve preconditions, and instruct their subordinates. Subordinates’ plans, by contrast, include awaiting instructions from the leader and passively absorbing any other related information that comes their way.
The leader’s script for collaboration is shown below. Since this is a meta-plan, it can be applied to many specific plans and starts with few details populated. Part of its task is to fill in task-specific details. Select comments are provided for this human-readable formalism. For further details on this formalism, please refer to [11, 12, 9].
@@
As the leader begins to process the meta-plan, it will select the domain plan to pursue. In the example below, the selected plan is SEARCH-FOR-LOST-OBJECT. The details of the selected plan can be placed into the [RUN-PLAN] section of COLLABORATIVE-ACTIVITY, and any preconditions that must be resolved can be added to the [PRECONDITIONS] section. An excerpt from the leader’s script for COLLABORATIVE-ACTIVITY is shown below.
@@
Once the leader has worked through the preconditions, it can move to SUGGEST-PLAN, in which it will share the domain plan with subordinates through natural language dialog. These agents will receive this information and adopt their assignment(s) within the suggested plan. This process results in the team adopting a shared collaborative plan.
@@ IV-B2 Low-level control
Behavior Trees (BTs) in the tactical layer for both robots follow a standardized design template prioritizing safety and operational needs as shown in Fig. 3.c. This design leverages left-to-right execution priority, ensuring the robot’s critical needs and priorities. Collision avoidance sub-trees are positioned leftmost, followed by robot-needs sub-trees, then action-command sub-trees. On the far right, fallback sub-trees handle behaviors like random exploration or returning to the base station. An excerpt from the UGV’s BT is shown in Fig. 3.d.
@@ V Evaluation Results
The simulation environment features an integrated chat UI through which the human operator (Danny) interacts with the robots. As shown in Fig. 4(1-5), the interface comprises multiple components: a central chat window, visualization widgets for the robots’ language interpretations and visual perceptions (VMRs), natural language representations of robot cognition, and structured displays of their goals and plans. This comprehensive design offers transparency into the robots’ decision-making processes and cognitive operations. The analysis below refers to specific labeled elements in Fig. 4. Please view the accompanying video for a better understanding. Additional information is also available on the HARMONIC project page.  [^1] https://sanjayovs.github.io/HARMONIC2025IROS/
@@ V-2 Information Gathering
The UGV proceeds to verify preconditions for the SEARCH-FOR-LOST-OBJECT plan (Fig. 4.3). Although the object type is already known, information about its features and last-seen location would be useful. The UGV queries Danny for this information (M2/M4), and receives responses (M3/M5). The latter response prompts the system to prioritize the entry-way in the search sequence (ii).
@@ VI Conclusion and Future Work
Our demonstration system of a simulated search task involving two robots and a human teammate illustrates the potential for HARMONIC to be applied to real-world applications. Future research will focus on expanding the range of tasks and scenarios in which HARMONIC agents can successfully operate; enhancing the robots’ learning capabilities; conducting real-world trials using physical robots to validate the framework’s effectiveness in diverse operational contexts; developing collaborative online planning strategies; and enhancing real-time adaptability in dynamic environments.
@@ ACKNOWLEDGMENT
This research was supported in part by grant ##N00014-23-1-2060 from the U.S. Office of Naval Research. Any opinions or findings expressed in this material are those of the authors and do not necessarily reflect the views of the Office of Naval Research.

### v3

@@ IV-B Development of Distributed Plans
Plans are distributed between strategic and tactical components in both the UGV and the drone. The high-level plans (ontological script instances) are divided into steps communicated to the tactical layer as action commands, as described earlier.
@@ IV-B1 High-level planning
The ontologies of agents in HARMONIC contain scripts for complex events that a particular agent can perform or understand other agents performing. Plans are instances of scripts with parameter values set for the situation at hand. These are instantiated by executing one or more of the ontologically stored meta-scripts, that is, scripts that HARMONIC robots instantiate to guide the generation of plans from scripts in their operational domain. One such meta-script is the COLLABORATIVE-ACTIVITY script, whose purpose is to help agents operating in teams organize themselves to accomplish a goal.
The details of the COLLABORATIVE-ACTIVITY script vary based on the agent’s role in the team. For example, the team (or task) leader’s script is typically more complex than the scripts of its subordinates since team leaders must select a domain script to instantiate into the plan to accomplish the target goal, resolve preconditions, and instruct their subordinates. Subordinates’ scripts, by contrast, include awaiting instructions from the leader and passively absorbing any other related information that comes their way.
The leader’s script for collaboration is shown below. Since this is a meta-script, it can be applied to many specific scripts and starts with few parameter values populated. Part of its task is to fill in task-specific parameter values. Select comments are provided for this human-readable formalism. For further details on this formalism, please refer to [11, 12, 9].
@@
Having instantiated an instance of COLLABORATIVE-ACTIVITY as a meta-plan, the leader will select the domain script to pursue and instantiate as a domain plan. In the example below, the selected script is SEARCH-FOR-LOST-OBJECT. The details of the plan instantiated from the selected script can be placed into the [RUN-PLAN] section of COLLABORATIVE-ACTIVITY, and any preconditions that must be resolved can be added to the [PRECONDITIONS] section. An excerpt from the leader’s script for COLLABORATIVE-ACTIVITY is shown below.
@@
Once the leader has worked through the preconditions, it can move to SUGGEST-PLAN, in which it will share the domain plan with subordinates through natural language dialog. These subordinates will receive this information and adopt their assignment(s) within the suggested plan, resulting in the team adopting a shared collaborative plan.
@@ IV-B2 Low-level control
For low-level planning and control on the tactical layer, we use Behavior Trees (BTs).
BTs provide a flexible and intuitive way to design control actions and define task-planning hierarchies for robots [13]. Their modular design enables easy modification and scaling of behaviors, making BTs particularly suitable for complex hierarchical task planning, in the tactical layer.
BTs for both robots follow a standardized design template prioritizing safety and operational needs as shown in Fig. 3.c. This design leverages the conventional left-to-right execution priority, ensuring the robot’s critical needs and priorities. Collision avoidance subtrees are positioned leftmost, followed by robot-needs sub-trees, then action-command sub-trees. On the far right, fallback sub-trees handle behaviors like random exploration. An excerpt from the UGV’s BT is shown in Fig. 3.d.
@@ V Evaluation Results
The simulation environment features an integrated chat UI through which the human operator (Danny) interacts with the robots. As shown in Fig. 4(1-5), the interface comprises multiple components: a central chat window, visualization widgets for the robots’ language interpretations and visual perceptions (VMRs), natural language representations of robot cognition, and structured displays of their goals and plans. This design offers transparency into the robots’ decision-making processes and cognitive operations. The analysis below refers to specific labeled elements in Fig. 4. Please view the accompanying video for a better understanding.
@@ V-2 Information Gathering
The UGV proceeds to verify preconditions for the SEARCH-FOR-LOST-OBJECT plan (Fig. 4.3). Although the object type is already known, information about its features and last-seen location would be useful. The UGV queries Danny for this information (M2/M4), and receives responses (M3/M5). The latter response prompts the system to prioritize entry-way in the search sequence (ii).
@@ VI Conclusion and Future Work
Our demonstration system of a simulated search task involving two robots and a human teammate illustrates the potential for HARMONIC to be applied to MRS problems. Future research will focus on expanding the range of tasks and scenarios in which HARMONIC agents can successfully operate; enhancing the robots’ learning capabilities; conducting real-world trials using physical robots to validate the framework’s effectiveness in diverse operational contexts; and developing collaborative online planning strategies in dynamic environments;
@@ ACKNOWLEDGMENT
This work was supported in part by ONR Grant #N00014-23-1-2060. The views expressed are those of the authors and do not necessarily reflect those of the Office of Naval Research.